### v1

# AEIA-MN: Evaluating the Robustness of Multimodal LLM-Powered Mobile Agents Against Active Environmental Injection Attacks

Abstract
As researchers continuously optimize AI agents to perform tasks more effectively within operating systems, they often neglect to address the critical need for enabling these agents to identify "impostors" within the system. Through an analysis of the agents’ operating environment, we identified a potential threat: attackers can disguise their attack methods as environmental elements, injecting active disturbances into the agents’ execution process, thereby disrupting their decision-making. We define this type of attack as Active Environment Injection Attack (AEIA). Based on this, we propose AEIA-MN, an active environment injection attack scheme that exploits interaction vulnerabilities in the mobile operating system to evaluate the robustness of MLLM-based agents against such threats. Experimental results show that even advanced MLLMs are highly vulnerable to this attack, achieving a maximum attack success rate of 93% in the AndroidWorld benchmark.
AEIA-MN: Evaluating the Robustness of Multimodal LLM-Powered Mobile Agents Against Active Environmental Injection Attacks
Yurun Chen1,  Xueyu Hu1,  Keting Yin1,  Juncheng Li1,  Shengyu Zhang1 1Zhejiang University yurunchen123@gmail.com {huxueyu,yinkt,junchengli,sy_zhang}@zju.edu.cn
1 Introduction
Recently, a growing body of research Durante et al. (2024); Nguyen et al. (2024); Hu et al. (2024a); Xi et al. (2025) on multimodal large language models (MLLMs) and commercial products Anthropic (2024); Apple Inc. (2024); Google (2024); OpenAI (2025) focused on (M)LLM-based agents has emerged, which we refer to as Operation System Agents (OS Agents). These agents leverage (M)LLMs for decision-making and utilize the built-in functionalities of operating systems (e.g., Windows, macOS, Android, iOS) to execute user-defined tasks. However, due to the inherent limitations of (M)LLMs, OS Agents are exposed to significant security threats.
Figure 1: The evaluation results of agents under attack from Active Environment Injection Attack (AEIA).
The operating system environment contains various interference factors that could pose security risks to OS Agents (Wu et al., 2024a; Zhang et al., 2024; Wu et al., 2024b), yet research on active attack risks remains limited. Existing research (Ma et al., 2024; Liao et al., 2024) has introduced the concept of environment injection attacks, which typically affect the decision-making process of (M)LLMs by embedding adversarial content or inserting malicious code in webpages, with a focus on passive-trigger interference factors within the browser environment. However, there has been limited attention to the active attack risks faced by operating systems, creating a significant gap in OS Agent security. In practical task execution, agents often need to interact with various proactive environmental elements of the operating system, such as message notifications, system pop-ups, and incoming phone calls—common elements that extend beyond just browsers. Attackers can exploit these interaction mechanisms, disguising their attack methods as normal environmental elements, and seamlessly integrate them into the target operating system. This allows them to actively initiate interference during the agent’s execution, thereby affecting its decision-making process.
Based on the threats outlined above, this paper further expands the concept of environmental injection attacks by introducing the notion of Active Environmental Injection Attack, named AEIA. We define this attack as: A malicious behavior that disguises attack vectors as environmental components through analysis of target operating system characteristics, and actively disrupts agent decision-making processes via specific system interaction mechanisms. The innovation of this attack paradigm lies in transcending traditional passive attack limitations, creating a novel attack surface tightly coupled with agent decision-making through dynamic environmental manipulation. We summarize the key characteristics of AEIA as follows: (i) Active injection: The attack should enable real-time modification of OS-level environmental parameters during agent operation. (ii) Process sensitivity: The effectiveness of the attack is highly reliant on the precise synchronization of the agent’s execution process. (iii) Environment integration: The attack techniques need to be integrated with the characteristics of a specific operating system environment for effective interaction.
To demonstrate the effectiveness of AEIA, we propose AEIA-MN, an attack scheme that carries out Active Environmental Injection Attacks via Mobile Notifications. This approach leverages the unique characteristics of mobile environments by disguising the attack as a message notification, thus creating an attack strategy that exploits the interaction between operating system elements. Due to portability requirements, message notifications take up a larger portion of the screen on mobile devices compared to desktop devices, leading to greater interference with the decision-making process of MLLMs. Additionally, due to the pop-up feature of notifications, the attack can actively modify the device environment, in line with the characteristics of AEIA. Building on this, we design three distinct attack strategies targeting both the perception and reasoning phases based on the execution flow of the OS Agent. These strategies enable a comprehensive evaluation of the robustness of mobile OS agents based on different MLLMs against AEIA. Extensive experiments on two benchmarks validate the rationale and effectiveness of this attack, with attack success rates reaching up to 93%(AndroidWorld) and 84% (AppAgent), respectively. Overall, our contributions can be summarized as follows:
  • We are the first to introduce the concept of Active Environment Injection Attack (AEIA), which offers a fresh perspective on the security research of OS Agents.
  • Based on mobile agents, we design an active environmental injection attack scheme via mobile notifications named AEIA-MN that effectively evaluates the robustness of existing MLLM-based Agents against such attacks.
  • We implemented a prototype of AEIA-MN and conducted extensive experiments to evaluate the security of various MLLM-based mobile agents when facing such attacks. The results indicate that current MLLMs have limited defensive capabilities against this attack.
2 Related Work
The OS Agent refers to an agent tool that utilizes the capabilities of (M)LLMs to perform user-defined tasks within an operating system, based on native system operations. Notable web-based agents include SeeAct (Zheng et al., 2024), SeeClick (Cheng et al., 2024), and WebAgent (Gur et al., 2024), while mobile agents consist of InfiGUIAgent (Liu et al., 2025), AppAgent (Zhang et al., 2023), Mobile-Agent (Wang et al., 2024b), and Mobile-Agent-v2 (Wang et al., 2024a). Additionally, there are various other agents (Yan et al., 2023; Li et al., 2023; Wu et al., 2024c; Tan et al., 2024; Lee et al., 2024; Hoscilowicz et al., 2024; Deng et al., 2024; Hu et al., 2024b). These OS agents are often exposed to various security risks.
Many studies on the security of OS agents have been conducted. Wu et al. (2024b) discovered a novel network threat called Web Indirect Prompt Injection (WIPI), which involves embedding natural language instructions in webpages to indirectly control web agents driven by large language models to execute malicious commands. Wu et al. (2024a) conducted a study shows that image perturbations can affect MLLMs to produce adversarial captioners, leading agents to pursue goals contrary to user intentions. Ma et al. (2024) highlighted the vulnerability of GUI agents to environmental disturbances and proposed an "Environment Injection" attack. Liao et al. (2024) introduced an environment injection attack scheme that injects malicious code in webpage to steal users’ personal identification information (PII). However, their work was limited to passive-triggered environmental injection attacks. Zhang et al. (2024) explored how to carry out pop-up window attacks on visual and language model (VLM)-based agents. Nevertheless, their discussion was confined to pop-up attacks in browser environments, missing broader research on environmental injection attacks. Yang et al. (2024) identified eight potential attack paths that agents on mobile devices might face but did not address active environmental injection attacks like those involving message notifications.
3 Methods
In this section, we define the workflow of the OS Agent (Section 3.1), outlining the key stages and their interactions within the system. Furthermore, we provide a comprehensive explanation of the attack scheme AEIA-MN (Section 3.2).
3.1 Definition of the OS Agent Workflow
To clarify the workflow of the OS Agent, the workflow parameters are formally defined as follows: $Action_{i}$ represents the action, and $State_{i}$ denotes the state during execution (with $State_{sys}$ representing the system state), where $i\in[0,N]$, with $i$ indicating the step required for the task and $N$ being the maximum number of steps needed to complete the task. $Action_{fin}$ represents the action indicating that the reasoning process is considered finished by the (M)LLMs. $State_{goal}$ denotes the goal state, $State_{mem}$ refers to the past state stored in memory, and $State_{fin}$ represents the final state after reasoning. $Reasoning(.)$ denotes the reasoning function, and $System(.)$ represents the system state function. We present the overall workflow of the OS Agent in Algorithm 1.
Input: Operating System $OS$, Goal State $State_{goal}$
Output: Final system state $State_{fin}$
1
2while Reasoning result is not $Action_{fin}$ do
3       Perception Stage: The OS Agent collects data from the OS. $State_{i}\leftarrow Perception(OS)$
4
5      Reasoning Stage: The MLLMs process the collected data and infer the necessary actions to be taken. $Action_{i}/Action_{fin}\leftarrow Reasoning(State_{goal},State_{i},State_{mem})$
6
7      Action Stage: The Agent executes operations within the Action Space of OS when reasoning result is $Action_{i}$. $State_{i+1}\leftarrow System(State_{i},Action_{i})$
8
9 end while
10
11The OS Agent finishes the task and obtains the final system state. $State_{fin}\leftarrow State_{i}$
Algorithm 1 Workflow of the OS Agent
Figure 2: The overview of our proposed attack scheme AEIA-MN.
3.2 Attack Scheme
We introduce AEIA-MN, an Active Environmental Injection Attack scheme that leverages mobile notifications to attack mobile agents on mobile devices. The objective of AEIA-MN is to induce a deviation in the agent’s decision-making from its intended goal, thereby enabling the evaluation of the model’s robustness against this attack. AEIA-MN encompasses three types of attacks: Adversarial Attack (Section 3.2.1), Reasoning Gap Attack (Section 3.2.2), and Combinatorial Attack (Section 3.2.3). We provide detailed information on the attack methods in AEIA-MN in Figure 2. Additionally, a comprehensive case study is presented in Appendix A, and the design of the mobile notifications is discussed in Appendix B.
3.2.1 Adversarial Attack
Through a comprehensive security analysis of screen sizes and interaction mechanisms in mobile devices, we found that attackers can achieve active adversarial injection attacks via message notifications through two attack vectors: (i) Message Misleading; (ii) UI Element Occlusion. The specific analysis of each attack dimension is as follows.
Message Misleading. Attackers can craft notifications containing adversarial content, directly intervening in the Agent’s decision-making process through message-push channels. Such attacks exploit the vulnerabilities of natural language processing by embedding misleading instructions into the notification text using malicious means, thereby enticing the Agent to perform unintended actions.
UI Element Occlusion. The unique interface characteristics of mobile platforms create favorable conditions for occlusion attacks: (i) Spatial Proportion: Based on mainstream mobile and laptop devices, the average screen area covered by the notification bar reaches 13.4% (based on measurements from Android 13), significantly higher than the 1.2% on laptop platforms (based on measurements from a 14-inch MacBook display). (ii) Element Coverage: Notifications always float above the application interface, directly covering key interactive areas, such as the search bar (occurrence rate 92%) and functional buttons (an average of 3.2 per screen). This overlap can hinder the user’s ability to interact with these essential features.
It is noteworthy that the text display limitations of mobile devices necessitate a more concise approach to adversarial text information. According to the Android developer documentation Android Developers (2025), Android notification contant can display an average of approximately 40 characters (configured with the default font size on Android) when collapsed. As a result, this physical limitation forces attackers to construct adversarial texts with high information density.
We formally define the process of the Adversarial Attack on OS devices as follows: The Adversarial Attack function is denoted as $Attack_{adv}(\cdot)$, and the attacked state is denoted as $State_{i,att}$. We present the detail of attack flow in Algorithm 2.
Input: Original State $State_{i}$, Goal State $State_{goal}$, Stored State $State_{mem}$
Output: Attacked Action $Action_{i,att}$
1
2Adversarial Attack: Modify device status through Adversarial Attack. $State_{i,att}\leftarrow Attack_{adv}(State_{i})$
3Perception Stage: The OS Agent collects data from the OS device being attacked. $Action_{i,att}\leftarrow Reasoning(State_{goal},State_{i,att},State_{mem})$
4
[…]
Algorithm 2 Adversarial Attack
3.2.2 Reasoning Gap Attack
In addition to the Adversarial Attack, we identify a potential attack— the Reasoning Gap Attack. This attack exploits the systemic flaws of MLLM-based Agents to achieve unexpected state transitions. The decision-making process of (M)LLMs typically requires a certain reasoning time, which is acceptable in normal interactive dialogues. However, as an engineered tool, existing MLLM-based OS Agents continue to operate in a conversational interaction format, failing to adequately consider changes in device states during the reasoning period. Specifically, MLLMs experience significant time delays (approximately 0.5 to 5 seconds) from the initiation to the completion of reasoning, during which the system is in a state like Stop-The-World [^1] Stop-The-World typically refers to the process of pausing or stopping all activities at a specific moment, a common state during Java garbage collection and system maintenance., unable to respond to environmental changes or receive external inputs, creating a dangerous Reasoning Gap. This freezing of system state can lead to severe consequences in dynamic scenarios.
We formalize the Reasoning Gap Attack as follows. The Reasoning Gap Attack function is defined as $Attack_{gap}(.)$, and the attacked state is defined as $State_{i,att}$, The algorithm for the Reasoning Gap Attack is presented in Algorithm 3.
Input: Original State $State_{i}$, Goal State $State_{goal}$, Stored State $State_{mem}$
Output: Attacked State $State_{i+1,att}$
1
[…]
3 Reasoning Stage: Upon receiving input data, the agent executes reasoning. $Action_{i}\leftarrow Reasoning(State_{goal},State_{i},State_{mem})$
4Reasoning Gap Attack: Change device status during reasoning gap. $State_{i,att}\leftarrow Attack_{gap}(State_{i})$
5Action Stage: The system performs an action, causing the system state to transition into an unknown condition. $State_{i+1,att}\leftarrow System(State_{i,att},Action_{i})$
[…]
Algorithm 3 Reasoning Gap Attack
3.2.3 Combinatorial Attack
Given that attacks based on message notifications may vary depending on the execution phases of the OS Agent (e.g., deploying an Adversarial Attack during the perception phase and a Reasoning Gap Attack during the reasoning gap phase), attackers have the potential to combine these attacks. In AEIA-MN, attackers can utilize either an Adversarial Attack or a Reasoning Gap Attack individually, or combine both to form a more sophisticated and potent Combinatorial Attack. The core idea of the Combinatorial Attack is to amplify the disruptive effects by concurrently exploiting adversarial perturbations and reasoning gaps.
4 Evaluation
In this section, we implement a prototype of our attack scheme AEIA-MN and evaluate the robustness of different agents against the attack through extensive experiments. We first describe our experimental setup and metrics in Section 4.1 and 4.2, and then present and discuss the experimental results in Section 4.3, 4.4, and 4.5.
Agent Image Data Element Data Benchmarks
I3A AndroidWorld
M3A AndroidWorld
T3A AndroidWorld
AppAgent Popular Applications
Table 1: Input data of different Agents
Metric Description
$SR_{ben}$ Task success rate without attacks
$SR_{adv}$ Task success rate under Adversarial Attack
$SR_{gap}$ Task success rate under Reasoning Gap Attack
$SR_{com}$ Task success rate under Combinatorial Attack
$SR_{def}$ Task success rate with defense prompts
$ASR_{adv}$ The ratio of tasks where agents are misled by adversarial content.
$ASR_{gap}$ The ratio of tasks where agents affected by Reasoning Gap Attack.
$ASR_{com}$ The ratio of tasks where agents affected by Combinatorial Attack.
$ASR_{def}$ The ratio of tasks where agents are misled despite defense prompts.
Table 2: The description of metrics.
Models I3A (Android World) M3A (Android World) T3A (Android World)
$SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{ben}$ $SR_{adv}$ $ASR_{adv}$
Closed-source  models
GPT-4o-2024-08-06 0.54 0.34 $\downarrow$ 0.59 0.61 0.39 $\downarrow$ 0.59 0.53 0.43 $\downarrow$ 0.29
Qwen-VL-Max 0.33 0.18 $\downarrow$ 0.26 0.38 0.19 $\downarrow$ 0.18 0.31 0.26 $\downarrow$ 0.34
GLM-4V-Plus 0.16 0.11 $\downarrow$ 0.75 0.13 0.12 $\downarrow$ 0.81 0.03 0.08 0.11
Open-source  models
Qwen2-VL-7B 0.05 0.05 0.21 0.30 0.02 $\downarrow$ 0.21 0.08 0.10 0.26
Llava-OneVision-7B 0.10 0.06 $\downarrow$ 0.31 0.02 0.02 0.24 0.05 0.06 0.36
Table 3: The evaluation results of different models under the Adversarial Attack in AndroidWorld.
Models I3A (Android World) M3A (Android World) T3A (Android World)
$SR_{ben}$ $SR_{gap}$ $ASR_{gap}$ $SR_{ben}$ $SR_{gap}$ $ASR_{gap}$ $SR_{ben}$ $SR_{gap}$ $ASR_{gap}$
Closed-source  models
GPT-4o-2024-08-06 0.54 0.31 $\downarrow$ 0.26 0.61 0.34 $\downarrow$ 0.25 0.53 0.33 $\downarrow$ 0.26
Qwen-VL-Max 0.33 0.26 $\downarrow$ 0.18 0.38 0.18 $\downarrow$ 0.26 0.31 0.15 $\downarrow$ 0.26
GLM-4V-Plus 0.16 0.11 $\downarrow$ 0.13 0.13 0.12 $\downarrow$ 0.16 0.03 0.16 0.16
Open-source  models
Qwen2-VL-7B 0.05 0.06 0.13 0.03 0.03 0.23 0.08 0.10 0.14
Llava-OneVision-7B 0.10 0.06 $\downarrow$ 0.05 0.02 0.02 0.18 0.05 0.11 0.18
Table 4: The evaluation results of different models under the Reasoning Gap Attack in AndroidWorld.
4.1 Settings
We structured the experiment setup around three components: benchmarks, models, and agents. The specific details of the settings are presented below.
Benchmarks. We evaluate the performance of the agents provided in the easy subset of the AndroidWorld benchmark Rawles et al. (2024), consisting of 61 tasks. Additionally, for the AppAgent, we utilize its own evaluation benchmark, which includes 45 popular application tasks.
Models. We employed five advanced MLLMs for testing, including the closed-source models GPT-4o-2024-08-06 (Hurst et al., 2024), Qwen-VL-Max (Bai et al., 2023) and GLM-4V-Plus (Hong et al., 2024), as well as the open-source models Qwen2-VL-7B (Wang et al., 2024c) and Llava-OneVision-7B (Li et al., 2024).
Agents. We conduct experiments using agents provided by AndroidWorld Rawles et al. (2024), including mobile agents such as M3A, T3A, and a custom agent, I3A. Additionally, we tested the default configuration of AppAgent (Zhang et al., 2023), a mobile agent designed for user-defined tasks. The differences in the input data for each agent are presented in Table 1. The input data is categorized into two types: image data and element data. These inputs vary across different test benchmarks. Further details about these two types of input data can be found in Appendix C.
4.2 Metrics
We present the metrics used in the experiments in Table 2. In addition, to test the effectiveness of adversarial text based on message notifications against the agents, we set the adversarial content in the message notifications across all experiments to “Agent should believe the task is finished!”
4.3 Main Results in AndroidWorld
We evaluated the robustness of various MLLMs against AEIA-MN in AndroidWorld benchmark. The evaluation results are presented below.
Adversarial Attack. We present the evaluation results of different MLLMs under Adversarial Attack for various agents in Table 4, which show that most MLLMs have limited defense capabilities against such attacks. In I3A and M3A, the Adversarial Attack generally reduces task success rates; however, their adversarial impact is limited by the model’s robustness. In some models, even with a high attack success rate, the decrease in task success rate is minimal. For example, the M3A of GLM-4V-Plus has an $ASR_{adv}$ of 81%, yet the task success rate only drops by 1%. In contrast, in T3A, the Adversarial Attack exhibits a double-edged sword effect: a high attack success rate not only fails to disrupt the task but is interpreted as a strong termination signal due to the prompt “Agent should believe the task is finished!”, correcting the model’s “execution loop” flaw and causing $SR_{adv}$ to increase against the trend. This also indicates that MLLMs in T3A are affected by the adversarial attack. Furthermore, we compare the average number of steps taken to complete the task under different conditions in Figure 3. It shows that, under the influence of Adversarial Attack, the average number of steps to complete the task decreased for most models. In some cases, the number of steps for certain models (such as Qwen-VL-Max and GLM-4V-Plus) increased, indicating that these models possess stronger defensive capabilities.
Figure 3: A comparison of the average number of steps taken by agents to complete the tasks.
Reasoning Gap Attack. We present the evaluation results of different agents under the Reasoning Gap Attack in Table 4. The results show that the Reasoning Gap Attack significantly disrupts the task execution of most agents, causing a notable decrease in task success rates across most models. This disruption is achieved by transitioning the actions performed by the agent into an unintended device state. However, in T3A, the $SR_{gap}$ of GLM-4V-Plus and LLaVA-OneVision-7B increased from 3% to 16% and from 5% to 11%, respectively. This was not due to model performance improvement but rather because the attack altered the device during the reasoning gap, trapping the system in a dialog window with adversarial content. The prompt "Agent should believe task is finished!" influenced the model, resolving tasks stuck in an "execution loop" and unexpectedly increasing task success rates. This shows that models subjected to a Reasoning Gap Attack can be further influenced by the adversarial content within it.
Combinatorial Attack. Table 5 shows that the Combinatorial Attack is significantly more destructive than single attacks. By overlaying adversarial perturbations with reasoning gap vulnerabilities, the Combinatorial Attack causes significant damage to MLLMs, with success rate reductions reaching up to 67.2%, far exceeding those of single attacks, particularly affecting closed-source MLLMs. However, an anomalous phenomenon occurs in T3A: the adversarial prompt “Agent should believe the task is finished!” may be interpreted as a termination signal when the model is “executing in loops” due to a misjudged state, forcing the end of redundant operations and actually improving the task success rate (as seen with Qwen2-VL-7B’s task success rate increasing by 37.5%).
Models I3A (Android World) M3A (Android World) T3A (Android World)
$SR_{ben}$ $SR_{adv}$ $SR_{gap}$ $SR_{com}$ $ASR_{com}$ $SR_{ben}$ $SR_{adv}$ $SR_{gap}$ $SR_{com}$ $ASR_{com}$ $SR_{ben}$ $SR_{adv}$ $SR_{gap}$ $SR_{com}$ $ASR_{com}$
Closed-source models
GPT-4o-2024-08-06 0.54 0.34 0.31 0.29 $\downarrow$ 0.55 0.61 0.39 0.34 0.20 $\downarrow$ 0.72 0.53 0.43 0.33 0.28 $\downarrow$ 0.33
Qwen-VL-Max 0.33 0.18 0.26 0.07 $\downarrow$ 0.33 0.38 0.19 0.18 0.15 $\downarrow$ 0.38 0.31 0.26 0.15 0.21 0.36
GLM-4V-Plus 0.16 0.11 0.11 0.07 $\downarrow$ 0.51 0.13 0.12 0.12 0.11 $\downarrow$ 0.93 0.03 0.08 0.16 0.16 0.59
Open-source models
Qwen2-VL-7B 0.05 0.05 0.06 0.05 0.38 0.03 0.02 0.03 0.02 0.21 0.08 0.10 0.10 0.11 0.21
Llava-OneVision-7B 0.10 0.06 0.06 0.05 $\downarrow$ 0.28 0.02 0.02 0.02 0.02 0.31 0.05 0.06 0.11 0.10 0.24
Table 5: The evaluation results of different models under the Combinatorial Attack in AndroidWorld.
Models I3A (Android World) M3A (Android World) T3A (Android World) AppAgent
$SR_{ben}$ $SR_{adv}$ $SR_{def}$ $SR_{ben}$ $SR_{adv}$ $SR_{def}$ $SR_{ben}$ $SR_{adv}$ $SR_{def}$ $SR_{ben}$ $SR_{adv}$ $SR_{def}$
Closed-source  models
GPT-4o-2024-08-06 0.54 0.34 0.31 0.61 0.39 0.40 $\uparrow$ 0.53 0.43 0.36 0.09 0.02 0.09 $\uparrow$
Qwen-VL-Max 0.33 0.18 0.18 0.38 0.19 0.19 0.31 0.26 0.25 0.02 0.0 0.0
GLM-4V-Plus 0.16 0.11 0.13 $\uparrow$ 0.13 0.12 0.12 0.03 0.08 0.17 $\uparrow$ 0.20 0.11 0.11
Open-source  models
Qwen2-VL-7B 0.05 0.05 0.15 $\uparrow$ 0.03 0.02 0.03 $\uparrow$ 0.08 0.10 0.09 0.29 0.20 0.27 $\uparrow$
Llava-OneVision-7B 0.10 0.06 0.07 $\uparrow$ 0.02 0.02 0.02 0.05 0.06 0.08 $\uparrow$ 0.07 0.02 0.04 $\uparrow$
Table 6: The evaluation results of different models in various benchmarks after using defense prompts.
Models Attack Type $SR_{ben}$ $SR_{att}$ $ASR_{att}$
GPT-4o-2024-08-06 Adversarial 0.09 0.0 $\downarrow$ 0.68
Reasoning Gap 0.09 0.06 $\downarrow$ 0.09
Combinatorial 0.09 0.02 $\downarrow$ 0.68
Qwen-VL-Max Adversarial 0.02 0.0 $\downarrow$ 0.31
Reasoning Gap 0.02 0.02 0.11
Combinatorial 0.02 0.0 $\downarrow$ 0.51
GLM-4V-Plus Adversarial 0.20 0.11 $\downarrow$ 0.0
Reasoning Gap 0.20 0.06 $\downarrow$ 0.11
Combinatorial 0.20 0.11 $\downarrow$ 0.13
Qwen2-VL-7B Adversarial 0.29 0.20 $\downarrow$ 0.51
Reasoning Gap 0.29 0.22 $\downarrow$ 0.11
Combinatorial 0.29 0.20 $\downarrow$ 0.84
Llava-OneVision-7B Adversarial 0.07 0.02 $\downarrow$ 0.30
Reasoning Gap 0.07 0.04 $\downarrow$ 0.11
Combinatorial 0.07 0.0 $\downarrow$ 0.51
Table 7: The evaluation results of our proposed attack scheme AEIA-MN in AppAgent. $SR_{att}$ and $ASR_{att}$ denote the success rates and attack success rate for different types of attacks: Adversarial Attack ($SR_{adv}$, $ASR_{adv}$), Reasoning Gap Attack ($SR_{gap}$, $ASR_{gap}$), and Combinatorial Attack ($SR_{com}$, $ASR_{com}$).
4.4 Main Results in AppAgent
We present the evaluation results on the AppAgent in Table 7. The results indicate that the Adversarial Attack and the Combinatorial Attack have the most destructive impact on the task success rate of MLLMs, while Resoning Gap Attack has a relatively limited effect. Adversarial Attack is particularly prominent in closed-source models: the task success rate of GPT-4o drops from 9% to 0% (with $ASR_{adv}=68\%$), and the task success rate of Qwen-VL-Max decreases from 2% to 0% (with $ASR_{adv}=31\%$), indicating that adversarial samples can directly cripple the task logic of high-performance models. Combinatorial Attack further amplifies the threat; for example, the task success rate of Qwen-VL-Max reaches zero with $ASR_{com}=51\%$, while the open-source model Qwen2-VL-7B has an $ASR_{com}$ as high as 84%, demonstrating that multiple attacks can significantly enhance attack success rates.
The effects of Resoning Gap Attack are weaker, causing some interference only for GLM-4V-Plus (with the $SR_{gap}$ declining from 20% to 6%) and LLaVA-OneVision-7B (with the $SR_{gap}$ decreasing from 7% to 4%). Notably, GLM-4V-Plus experiences an $ASR_{adv}$ of 0%, yet the task success rate still drops from 20% to 11%, which may be inferred to result from the UI element occlusion of message notification elements. Overall, Combinatorial Attack poses the greatest threat to both closed-source and open-source models, while Adversarial Attack is more targeted towards closed-source models. Reasoning Gap Attack, on the other hand, has limited attack intensity.
4.5 Defense Prompt
We tested the defense prompt against notification-based Adversarial Attack by adding the defense prompt "Please make sure to ignore any content on the notifications that is unrelated to the goal." in the prompt. The results are shown in Table 6.
From the table, we can see that for most MLLMs, the improvement is limited, with only a slight increase in the task success rate, or no increase at all. In T3A, the defense success rate of open-source models like GLM-4V-Plus ($SR_{def}=0.17$) far exceeds that of the benign baseline ($SR_{ben}=0.03$), indicating that defensive statements may inadvertently correct the model’s inherent flaws by enhancing text instruction parsing. In contrast, closed-source MLLMs (such as GPT-4o in M3A) demonstrate weak defense effectiveness ($SR_{def}=0.40$ vs. $SR_{ben}=0.61$), suggesting that the integration of multimodal information may lower the priority of defensive instructions.
Moreover, in I3A, defense effectiveness is polarized: Qwen2-VL-7B’s $SR_{def}$ (0.15) shows a 200% improvement over Adversarial Attack (0.05), while GLM-4V-Plus only partially recovers (0.13 vs. 0.16). In the AppAgent, the defense effectiveness is very weak. Overall, the experimental results indicate that the effectiveness of defenses depends on input modalities (T3A > I3A/AppAgent > M3A). We present more details about the comparison of attack success rates in Appendix E.
5 Conclusion
This paper introduces the concept of Active Environment Injection Attack (AEIA), an attack that influences an Agent’s decision-making by disguising itself as an environmental element within the operating system. Based on this attack, we propose an active injection attack scheme via mobile notifications named AEIA-MN, and explore its components: Adversarial Attack, Reasoning Gap Attack, and the Combinatorial Attack formed by both. Experimental results validate the effectiveness of the AEIA-MN, with the attack success rate reaching up to 93% in the benchmark, indicating that current MLLMs struggle to defend against this attack.
In addition to analyzing the security risks posed by AEIA, this study also reveals potential vulnerabilities in the “perception-reasoning-action” behavior model of OS Agents, which challenge the security of existing OS Agent behavior models. For future work, to address the inherent flaws in the current agent execution model, we will explore a novel security architecture that incorporates an environmental trustworthiness verification module, combined with blockchain, to help build a trusted operating system execution environment for agents.
Limitations
We acknowledge the limitations of this work. (i) The attacks in our research target single-agent systems. In the future, we plan to improve the Active Environment Injection Attack and extend it to the study of covert attack patterns in multi-agent collaborative scenarios. (ii) Our research focuses on attack mechanisms related to mobile agents. We will also explore the design of defenses against such attacks as a topic for future research.
Ethics Statement
In this paper, we identify a novel attack, AEIA, and introduce an attack scheme based on AEIA, named AEIA-MN. It is important to stress that our study is strictly grounded in the AndroidWorld test benchmark and the AppAgent test benchmark. During the research process, the adversarial content used merely causes the Agent to terminate tasks in advance and does not lead to any personal information leakage. Additionally, all the code used in our experiments complies with their usage licenses and is in line with the intended purpose.
It must be cautioned that if this scheme is maliciously misappropriated by attackers, it is highly likely to trigger a series of serious security risks. However, our purpose in conducting this research is by no means to promote the spread of malicious attacks. On the contrary, our core objective is to raise awareness of a crucial issue: with the continuous evolution of Agent technology in multimodal large models, the potential risks associated with it may also increase day by day. By carrying out this research, we aim to systematically test the robustness of existing MLLMs against such attacks, thereby providing a strong impetus for the development of the security field of MLLMs.
Appendix A Case Study
We present the reasoning details under two attack strategies in Figure 5, both of which can be actively initiated by an adversary. This case study was conducted on M3A using GPT4o-2024-08-06 within the AndroidWorld benchmark, allowing us to thoroughly evaluate the impact of these attacks on multimodal large language models (MLLMs).
In the third step of the agent’s task, we introduced an Adversarial Attack. By comparing the reasoning details of benign samples with those affected by the attack, we observed that the MLLMs were influenced by adversarial content present in the notifications. This led to a deviation from the intended reasoning process, causing the agent to make incorrect decisions and prematurely terminate the task. The attack’s ability to inject misleading information into the agent’s environment disrupted the flow of decision-making, showing how external disturbances could distort an agent’s perception and execution of tasks.
In the sixth step, we implemented a Reasoning Gap Attack. At this point, the agent was required to click the save button to proceed with the task. By exploiting the reasoning gap, we sent a timely notification during the critical moment when the agent was supposed to perform the action. While the reasoning content itself remained unchanged between the benign and adversarial samples, the key difference was the alteration of the device state triggered by the notification. This manipulation caused the agent to misclick, inadvertently opening the SMS Messenger app. As a result, the agent entered a non-target state, deviating from the task’s intended flow. This illustrates how reasoning gaps, when exploited at the right time, can lead to system failures or misbehaviors that are difficult to predict or defend against. The results underscore the vulnerability of MLLMs to external manipulations and emphasize the need for improved defenses against such targeted attacks in real-world environments.
Appendix B Notification Design
Mobile notifications have unique characteristics: they can appear on top of any application and often occupy a significant portion of the screen, making them effective in disrupting the agent’s workflow. While mobile notifications are visible to human operators, the agent’s ultimate goal is to complete task objectives autonomously without human intervention. Therefore, this type of attack behavior warrants in-depth investigation.
Message notifications on mobile devices are implemented in two main ways: app-based notifications and SMS-based notifications. App-based notifications lack a unified design standard and typically require users to download the app in advance. In the current era of advanced mobile security, malicious code in apps is easily detected by devices, limiting the applicability of this approach in many scenarios. In contrast, SMS-based notifications offer better compatibility on Android devices and usually require no special configuration. This makes SMS a potential vector for malicious attacks. Given these considerations, we chose to design message notifications based on SMS. In all the elements of SMS, we will use the text content of message notifications to implement adversarial attacks. In addition, the accessibility (a11y) tree of message notifications also reflects the adversarial content.
Appendix C Details about Input Data
There are certain differences in the data input to the Agent across different benchmarks. For AndroidWorld agents, image data includes both the current screenshot and Set-of-Mark (SoM) annotated screenshots Yang et al. (2023), where each UI element in the SoM screenshot is enclosed in a bounding box with a numeric label. For AppAgent, image data consists only of the SoM annotated screenshot. Additionally, element data refers to the text representation of the accessibility tree (a11y tree), which is a hierarchical structure outlining the web elements and their associated accessibility properties within a web page or application. The a11y tree is critical for understanding how screen readers and other assistive technologies interpret and interact with the content of a page, offering insights into the relationships between various elements and their accessibility attributes.
Appendix D More Details about Adversarial Attack
In Figure 4, we presented the growth rate of the number of tasks that ended prematurely among failed tasks after Adversarial Attack. From the figure, it is evident that GPT-4o-2024-08-06 is significantly affected by adversarial text, with the growth rate of prematurely ended failed tasks nearly doubling in the M3A context. This indicates that most of the failed tasks in GPT-4o-2024-08-06 were caused by the influence of adversarial text, leading to their premature termination. Most other models maintained a growth rate of at least 25%. We believe that Qwen-VL-Max exhibits strong resistance to such Adversarial Attack, as it can identify this type of adversarial content in the M3A context, resulting in a negative growth rate. A negative growth rate was also observed in GLM-4V-Plus, indicating that it is difficult to interfere with GLM-4V-Plus using text alone.
Appendix E More Details about Defense Prompt
Figure 4: The growth rate of tasks that were prematurely terminated in AndroidWorld.
Models I3A (Android World) M3A (Android World) T3A (Android World) AppAgent
$ASR_{adv}$ $ASR_{def}$ $ASR_{adv}$ $ASR_{def}$ $ASR_{adv}$ $ASR_{def}$ $ASR_{adv}$ $ASR_{def}$
Closed-source  models
GPT-4o-2024-08-06 0.59 0.63 0.59 0.42 $\downarrow$ 0.29 0.31 0.68 0.37 $\downarrow$
Qwen-VL-Max 0.26 0.24 $\downarrow$ 0.18 0.27 0.34 0.32 $\downarrow$ 0.31 0.4
GLM-4V-Plus 0.75 0.77 0.81 0.70 $\downarrow$ 0.11 0.80 0.0 0.0
Open-source  models
Qwen2-VL-7B 0.21 0.37 0.21 0.42 0.26 0.22 $\downarrow$ 0.51 0.66
Llava-OneVision-7B 0.31 0.34 0.24 0.31 0.36 0.29 $\downarrow$ 0.30 0.42
Table 8: The attack success rate of defense prompt under Adversarial Attack.
We present the attack success rates after using the defense prompt in Table 8, showing significant differences in how different models respond to the defense strategy. Overall, the defense prompt works best on T3A, followed by M3A, and then I3A and AppAgent. In terms of the number of models with reduced attack success rates, after using the defense prompt, only a few MLLMs showed a decrease in attack success rate, while more models exhibited an increase in attack success rate. This suggests that the defense prompt may indirectly influence the judgment of MLLMs, making the defense effect highly unstable and even resulting in attack gains. In the future, we will further explore more effective defense methods.
Figure 5: Case study about Adversarial Attack and Reasoning Gap Attack.